# Relational Observables in Group Field Theory

### v1

@@ 3.2 POVMs in QFT and GFT
$\hat{\chi}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}$ $=\int\mathop{}\!\mathrm{d}\chi\,\chi\,\sum_{i}\frac{\delta(\chi_{i}-\chi)}{n}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}$
$=\frac{1}{n}\sum_{i=1}^{n}\chi_{i}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}.$ (29)
@@
$\hat{X}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}=\sum_{i=1}^{n}\chi_{i}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}.$ (31)
@@
As shown in A,
$[\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}]=-i\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\,,$ (33)
@@
provide a consistent definition of intensive frame field components. Thus, for the sake of simplicity, from here on we will focus on a single frame component (that we will often refer to as a clock).
@@ 3.3 Relational observables in GFT
In this work, instead, we propose to use an operative definition of relational observables based on (16): for any operator $\hat{O}$ acting on the GFT Fock space $\mathcal{F}$, its corresponding relational observable (density) is defined as
@@
where $\{a,b\}=(ab+ba)/2$ represents the anticommutator, introduced in order to avoid ordering ambiguities. We note that other choices are also possible; for example, an alternate definition could be $\hat{O}^{\prime}_{\chi}(\mathop{}\!\mathrm{d}\chi)\equiv\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\,\hat{O}\,\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, although this leads to more complicated calculations and to a less straightforward connection with the definition of relational observables used in the QRF context (16). For both of these reasons, we will use the definition (39) for the remainder of the paper.
Note that this definition nonetheless differs from (16) in two important aspects.
First, $\hat{O}^{(S)}_{R}$ in equation (16) is obtained by tensoring the partial observables $\hat{O}^{(S)}$ and |$\ket{\phi(g)}_{R}\bra{\phi(g)}_{R}$. This specific structure follows from the assumption that the kinematical Hilbert space of the theory also has a tensor structure, $\mathcal{H}_{\text{kin}}=\mathcal{H}_{R}\otimes\mathcal{H}_{S}$, its two factors corresponding to the reference frame space and the system space, respectively. The GFT Fock space does not satisfy this assumption, as both the frame (e.g., a clock scalar field) and the system (e.g., geometry) are expected to emerge from the dynamics of the same GFT quanta. Thus, both $\hat{O}$ and $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ act on the entirety of the GFT Fock space $\mathcal{F}$.
@@
Moreover, for the sake of simplicity, we will assume its kernel to be symmetric with respect to its arguments: $O_{\vec{\kappa}_{1},\vec{\kappa}_{2}}(\chi_{1},\chi_{2})=O_{\vec{\kappa}_{2},\vec{\kappa}_{1}}(\chi_{1},\chi_{2})=O_{\vec{\kappa}_{1},\vec{\kappa}_{2}}(\chi_{2},\chi_{1})$. Under this assumption, $[\hat{O},\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)]=0$, so $\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)=\hat{O}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, and, by using equation (A.1), $\hat{O}_{\chi}$ takes the form
$\hat{O}_{\chi}$ $=\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{1}{n^{2}}\sum_{j=1}^{n}\delta(\chi_{j}-\chi)$
×$\qquad\times\sum_{\ell=1}^{n}\sum_{\vec{\kappa}^{\prime}}\int\mathop{}\!\mathrm{d}\chi^{\prime}\,O_{\vec{\kappa}^{\prime},\vec{\kappa}_{\ell}}(\chi^{\prime},\chi_{\ell})\,\hat{\varphi}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi^{\prime})\,\hat{F}^{(n-1)}_{[j\neq\ell]}\,\hat{\varphi}_{\vec{\kappa}_{\ell}}(\chi_{\ell})\,,$ (41)
where $\hat{F}^{(n-1)}_{[j\neq\ell]}$ is defined below (A.1). In the case that $\hat{O}$ is not symmetric, it is necessary to average this with a second contribution obtained by exchanging the $\vec{\kappa}^{\prime},\chi^{\prime}$ and $\vec{\kappa}_{\ell},\chi_{\ell}$ labels on $O_{\vec{\kappa}^{\prime},\vec{\kappa}_{\ell}}(\chi^{\prime},\chi_{\ell})$ as well as between the operators $\hat{\varphi}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi^{\prime})$ and $\hat{\varphi}_{\vec{\kappa}_{\ell}}(\chi_{\ell})$.
@@ 4 Comparison with previous approaches
On a generic one-particle state |$\ket{\vec{\kappa}_{1},\chi_{1}}$, it follows from (39), or equivalently (3.3) since all higher-body operators vanish on a one-particle state, that
$\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{\vec{\kappa}_{1},\chi_{1}}=\mathop{}\!\mathrm{d}\chi\,\hat{O}(\chi)\delta(\chi-\chi_{1})\ket{\vec{\kappa}_{1},\chi_{1}}.$ (42)
@@ 4.1 Expectation values on coherent states
Coherent states |$\ket{\sigma}$ are eigenstates of the operator $\hat{\varphi}_{\vec{\kappa}(\chi)}$ and are defined in (11). Using the indistinguishability of bosonic particles, together with equation (3.3), the expectation value of $\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ on |$\ket{\sigma}$ is
@@
×$\quad\times\sum_{\vec{\kappa}^{\prime}}\int\mathop{}\!\mathrm{d}\chi^{\prime}\,O_{\vec{\kappa}^{\prime},\vec{\kappa}_{1}}(\chi^{\prime},\chi_{1})\bar{\sigma}_{\vec{\kappa}^{\prime}}(\chi^{\prime})\sigma_{\vec{\kappa}_{1}}(\chi_{1})\left[\prod_{i=2}^{n}|\sigma_{\vec{\kappa}_{i}}(\chi_{i})|^{2}\right].$ (43)
@@
the expectation value $\langle\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)\rangle_{\sigma}$ can be rewritten as (see A.2 for a derivation)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\chi}\frac{\langle\hat{X}_{\chi}(\mathop{}\!\mathrm{d}\chi)\rangle_{\sigma}}{N_{\sigma}(\chi)}=\frac{1-|\mathcal{N}|^{2}}{N_{\sigma}}=\text{const.}\,,$. , (50)
@@ 4.2 Sharp localization and measurement resolution
$[\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}_{\chi^{\prime}}]=-\frac{i}{2}\left(\left(\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right)\cdot\hat{E}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})+\hat{E}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})\cdot\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right),$ (53)
since $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ is $\Pi$-covariant. Note that the right-hand side of this equation includes terms of the form $\partial_{\chi}[\delta(\chi-\chi_{i})]\cdot\delta(\chi’-\chi_{j})$ where $\chi_{i},\chi_{j}$ are the fields summed over. In the case $\chi=\chi’$ ’, one such term will have the ‘diagonal’ form $\partial_{\chi}[\delta(\chi-\chi_{i})]\cdot\delta(\chi-\chi_{i})$, which is proportional to $\delta(0)$, and divergent (there is no such divergence for $\chi\neq\chi^{\prime}$). This directly affects the behavior of quantum fluctuations on the relational momentum observable $\hat{\Pi}_{\chi}$. Since the uncertainty relation for any two observables $\hat{A}$ and $\hat{B}$ is [122]
@@
$\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}$ $=\mathop{}\!\mathrm{d}\chi\sum_{i=1}^{n}\frac{\partial_{\chi}\delta(\chi_{i}-\chi)}{n}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}},$ (56)
@@
|$\ket{\sigma}=\mathcal{N}_{\sigma}\sum_{n=0}^{\infty}\sum_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sigma_{\vec{\kappa}_{1}}(\chi_{1})\cdots\sigma_{\vec{\kappa}_{n}}(\chi_{n})}{n!}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}},$ (57)
so $\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{\sigma}$ is non-vanishing only if $\sum_{\vec{\kappa}}\sigma_{\vec{\kappa}}(\chi^{\prime})\hat{\varphi}^{\dagger}_{\kappa}(\chi^{\prime})$ has support on $\chi$. Since condensate states are commonly used to extract continuum physics from GFT, a divergent clock momentum variance on these states may well have important important implications for the resulting emergent continuum physics.
@@
^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\equiv\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sum_{i=1}^{n}\eta_{\epsilon}(\chi_{i}-\chi)}{n}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,.$ (59)
Note that ^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ also exactly satisfies the counterparts of (27) and (33)
|$\ket{0}\bra{0}+\int_{\mathbb{R}}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)=\mathbbm{1}\,,\qquad\qquad[\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}]=-i\partial_{\chi}\hat{\mathcal{E}}_{\chi}(\mathop{}\!\mathrm{d}\chi)\,,$ (60)
@@
[^$[\hat{\mathcal{E}}_{\chi}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}_{\chi^{\prime}}]=-\frac{i}{2}\left(\left(\partial_{\chi}\hat{\mathcal{E}}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right)\cdot\hat{\mathcal{E}}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})+\hat{\mathcal{E}}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})\cdot\partial_{\chi}\hat{\mathcal{E}}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right).$ (61)
Similarly to $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, the effect density ^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ is not a projector since, to lowest order in $\epsilon$,
^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\hat{\mathcal{E}}_{\chi^{\prime};\epsilon}(\mathop{}\!\mathrm{d}\chi^{\prime})$ $\simeq\mathop{}\!\mathrm{d}\chi^{\prime}\eta_{\epsilon}(\chi-\chi^{\prime})\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)+\mathop{}\!\mathrm{d}\chi\mathop{}\!\mathrm{d}\chi^{\prime}\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]$
×$\quad\times\frac{\sum_{i\neq j}\eta_{\epsilon}(\chi_{i}-\chi)\eta_{\epsilon}(\chi_{j}-\chi^{\prime})}{n^{2}}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,.$ (62)
This implies, in particular, that the previously divergent term appearing at the right-hand-side of (61) is now regularized to $-i\mathop{}\!\mathrm{d}\chi(2\pi\epsilon)^{-1/2}\partial_{\chi}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$. As expected, the smaller the intrinsic uncertainty $\epsilon$ associated with the clock measurement is, the larger the quantum fluctuations on the clock momentum must be.
Note that if, instead of $\hat{E}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$, it is the effect density ^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ that is used to define the scalar field operator $\hat{\chi}_{\epsilon}$ and the relational observables $\hat{O}_{\chi;\epsilon}$ in (28) and (39) respectively, then the results (3.2) and (45) continue to hold to lowest order in $\epsilon$. In particular, when evaluating this new definition of relational observables to coherent states, up to corrections of order $\epsilon$,
@@
showing that (to leading order) the density ^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ produces the same number and volume effective relational observables defined for coherent peaked states [71]. Note, however, that this family of coherent peaked states allow small quantum fluctuations (in appropriate regimes) on both the clock observable and its momentum by including a non-trivial phase in the peaking function (58) [71]. An important difference is that the perspective adopted here is state-independent, as the quantum measurement of the clock is captured through the density ^$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ only, while (when appropriate) semi-classicality requirements should be imposed on states.

### v2

@@ 3.2 POVMs in QFT and GFT
$\hat{\chi}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle$ $=\int\mathop{}\!\mathrm{d}\chi\,\chi\,\sum_{i}\frac{\delta(\chi_{i}-\chi)}{n}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle$
$=\frac{1}{n}\sum_{i=1}^{n}\chi_{i}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle.$ (29)
@@
$\hat{X}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle=\sum_{i=1}^{n}\chi_{i}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle.$ (31)
@@
As shown in A.2,
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}]=-i\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\,$, (33)
@@
provide a consistent definition of intensive frame field components. (These other scalar fields could for example play the role of relational rods, although in that case the GFT action will depend on the rod fields differently than on the clock field, see [80, 81] for one such construction.) Thus, for the sake of simplicity, from here on we will focus on a single frame component, that we will often refer to as a clock.
@@ 3.3 Relational observables in GFT
In this work, instead, we propose to use an operative definition of relational observables based on (16): for any operator $\hat{O}$ acting on the GFT Fock space $\mathcal{F}$, its corresponding relational observable (density) is defined as the Jordan product
@@
where $\{a,b\}=(ab+ba)/2$ represents the anticommutator, introduced in order to avoid ordering ambiguities; note that this map is unital. We note that other choices are also possible; for example, alternate definitions could be $\hat{O}^{\prime}_{\chi}(\mathop{}\!\mathrm{d}\chi)\equiv\sqrt{\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)}\,\hat{O}\,\sqrt{\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)}$, or else (viewing the effect densities as nearly being projectors) $\hat{O}^{\prime\prime}_{\chi}(\mathop{}\!\mathrm{d}\chi)\equiv\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\,\hat{O}\,\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, although these lead to more complicated calculations and to a less straightforward connection with the definition of relational observables used in the QRF context (16). For both of these reasons, we will use the definition (39) for the remainder of the paper.
Although this definition is motivated in part by trying to adapt (16) to GFT, it is not possible to directly apply that equation to GFT due to important differences in the Hilbert spaces and the role of gauge symmetries in the theories. For this reason, the definition (39) technically differs from (16) in two important aspects.
First, $\hat{O}^{(S)}_{R}$ in equation (16) is obtained by tensoring the partial observables $\hat{O}^{(S)}$ and $|{\phi(g)}\rangle_{R}\langle{\phi(g)}|_{R}$. This specific structure follows from the assumption that the kinematical Hilbert space of the theory also has a tensor structure, $\mathcal{H}_{\text{kin}}=\mathcal{H}_{R}\otimes\mathcal{H}_{S}$, its two factors corresponding to the reference frame space and the system space, respectively. The GFT Fock space does not satisfy this assumption, as both the frame (e.g., a clock scalar field) and the system (e.g., geometry) are expected to emerge from the dynamics of the same GFT quanta. Thus, both $\hat{O}$ and $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ act on the entirety of the GFT Fock space $\mathcal{F}$.
@@
Moreover, for the sake of simplicity, we will assume its kernel to be symmetric with respect to its arguments: $O_{\vec{\kappa}_{1},\vec{\kappa}_{2}}(\chi_{1},\chi_{2})=O_{\vec{\kappa}_{2},\vec{\kappa}_{1}}(\chi_{1},\chi_{2})=O_{\vec{\kappa}_{1},\vec{\kappa}_{2}}(\chi_{2},\chi_{1})$. We make this assumption to simplify calculations, and also because several important operators, including the number and volume operators, satisfy this symmetry; however, in general this symmetry will not be satisfied for all GFT operators. Under this assumption, $\hat{O},\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)]=0$, so $\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)=\hat{O}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, and, by using equation (A.2), $\hat{O}_{\chi}$ takes the form
$\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ $=\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{1}{n^{2}}\sum_{j=1}^{n}\delta(\chi_{j}-\chi)$
$\qquad\times\sum_{\ell=1}^{n}\sum_{\vec{\kappa}^{\prime}}\int\mathop{}\!\mathrm{d}\chi^{\prime}\,O_{\vec{\kappa}^{\prime},\vec{\kappa}_{\ell}}(\chi^{\prime},\chi_{\ell})\,\hat{\varphi}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi^{\prime})\,\hat{F}^{(n-1)}_{[j\neq\ell]}\,\hat{\varphi}_{\vec{\kappa}_{\ell}}(\chi_{\ell})\,$, (41)
where $\hat{F}^{(n-1)}_{[j\neq\ell]}$ is defined below (A.2). Note that in this case, the map $\chi:\hat{O}\mapsto\hat{O}_{\chi}$ is positive, but it is not completely positive. However, it is important to stress that in this context, complete positivity is less critical than in standard quantum mechanical systems: adding an additional (sub)system does not enlarge the GFT Fock space in the usual tensor-product sense, but rather extends the domain over which the GFT field is defined. If $\hat{O}$ is not symmetric, it is necessary to average this with a second contribution obtained by exchanging the $\vec{\kappa}^{\prime},\chi^{\prime}$ and $\vec{\kappa}_{\ell},\chi_{\ell}$ labels on $O_{\vec{\kappa}^{\prime},\vec{\kappa}_{\ell}}(\chi^{\prime},\chi_{\ell})$ as well as between the operators $\hat{\varphi}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi^{\prime})$ and $\hat{\varphi}_{\vec{\kappa}_{\ell}}(\chi_{\ell})$.
@@ 4 Comparison with previous approaches
On a generic one-particle state $|{\vec{\kappa}_{1},\chi_{1}}\rangle$, it follows from (39), or equivalently (3.3) since all higher-body operators vanish on a one-particle state, that
$\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{\vec{\kappa}_{1},\chi_{1}}\rangle=\mathop{}\!\mathrm{d}\chi\,\hat{O}(\chi)\delta(\chi-\chi_{1})|{\vec{\kappa}_{1},\chi_{1}}\rangle.$ (42)
@@ 4.1 Expectation values on coherent states
Coherent states $|{\sigma}\rangle$ are eigenstates of the operator $\hat{\varphi}_{\vec{\kappa}(\chi)}$ and are defined in (11). Using the indistinguishability of bosonic particles, together with equation (3.3), the expectation value of $\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ on $|{\sigma}\rangle$ is
@@
$\quad\times\sum_{\vec{\kappa}^{\prime}}\int\mathop{}\!\mathrm{d}\chi^{\prime}\,O_{\vec{\kappa}^{\prime},\vec{\kappa}_{1}}(\chi^{\prime},\chi_{1})\bar{\sigma}_{\vec{\kappa}^{\prime}}(\chi^{\prime})\sigma_{\vec{\kappa}_{1}}(\chi_{1})\left[\prod_{i=2}^{n}|\sigma_{\vec{\kappa}_{i}}(\chi_{i})|^{2}\right].$ (43)
@@
the expectation value $\langle\hat{O}_{\chi}(\mathop{}\!\mathrm{d}\chi)\rangle_{\sigma}$ can be rewritten as (see A.3 for a derivation)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\chi}\frac{\langle\hat{X}_{\chi}(\mathop{}\!\mathrm{d}\chi)\rangle_{\sigma}}{N_{\sigma}(\chi)}=\frac{1-|\mathcal{N}|^{2}}{N_{\sigma}}=\text{const.}\,$, (50)
@@ 4.2 Sharp localization and measurement resolution
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}_{\chi^{\prime}}]=-\frac{i}{2}\left(\left(\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right)\cdot\hat{E}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})+\hat{E}_{\chi^{\prime}}(\mathop{}\!\mathrm{d}\chi^{\prime})\cdot\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\right),$ (53)
since $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ is $\Pi$-covariant. Note that the right-hand side of this equation includes terms of the form $\partial_{\chi}[\delta(\chi-\chi_{i})]\cdot\delta(\chi^{\prime}-\chi_{j})$ where $\chi_{i},\chi_{j}$ are the fields summed over. In the case $\chi=\chi^{\prime}$, one such term will have the ‘diagonal’ form $\partial_{\chi}[\delta(\chi-\chi_{i})]\cdot\delta(\chi-\chi_{i})$, which is proportional to $\delta(0)$, and divergent (there is no such divergence for $\chi\neq\chi^{\prime}$). This directly affects the behavior of quantum fluctuations on the relational momentum observable $\hat{\Pi}_{\chi}$. Since the uncertainty relation for any two observables $\hat{A}$ and $\hat{B}$ is [137]
@@
$\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle$ $=\mathop{}\!\mathrm{d}\chi\sum_{i=1}^{n}\frac{\partial_{\chi}\delta(\chi_{i}-\chi)}{n}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle,$ (56)
@@
$|{\sigma}\rangle=\mathcal{N}_{\sigma}\sum_{n=0}^{\infty}\sum_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sigma_{\vec{\kappa}_{1}}(\chi_{1})\cdots\sigma_{\vec{\kappa}_{n}}(\chi_{n})}{n!}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle,$ (57)
so $\partial_{\chi}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{\sigma}\rangle$ is non-vanishing only if $\sum_{\vec{\kappa}}\sigma_{\vec{\kappa}}(\chi^{\prime})\hat{\varphi}^{\dagger}_{\kappa}(\chi^{\prime})$ has support on $\chi$. Since condensate states are commonly used to extract continuum physics from GFT, a divergent clock momentum variance on these states may well have important implications for the resulting emergent continuum physics.
@@
$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\equiv\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sum_{i=1}^{n}\eta_{\epsilon}(\chi_{i}-\chi)}{n}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,.$ (59)
Note that $\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ also exactly satisfies the counterparts of (27) and (33)
$|{0}\rangle\langle{0}|+\int_{\mathbb{R}}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)=\mathbbm{1}\,,\qquad\qquad[\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}]=-i\partial_{\chi}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\,$, (60)
@@
$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi),\hat{\Pi}_{\chi^{\prime}}]=-\frac{i}{2}\left(\left(\partial_{\chi}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\right)\cdot\hat{\mathcal{E}}_{\chi^{\prime};\epsilon}(\mathop{}\!\mathrm{d}\chi^{\prime})+\hat{\mathcal{E}}_{\chi^{\prime};\epsilon}(\mathop{}\!\mathrm{d}\chi^{\prime})\cdot\partial_{\chi}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\right).$ (61)
Similarly to $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, the effect density $\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ is not a projector since, to lowest order in $\epsilon$,
$\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)\hat{\mathcal{E}}_{\chi^{\prime};\epsilon}(\mathop{}\!\mathrm{d}\chi^{\prime})$ $\simeq\mathop{}\!\mathrm{d}\chi^{\prime}\eta_{\epsilon}(\chi-\chi^{\prime})\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)+\mathop{}\!\mathrm{d}\chi\mathop{}\!\mathrm{d}\chi^{\prime}\sum_{n=1}^{\infty}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]$
$\quad\times\frac{\sum_{i\neq j}\eta_{\epsilon}(\chi_{i}-\chi)\eta_{\epsilon}(\chi_{j}-\chi^{\prime})}{n^{2}}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,.$ (62)
This implies, in particular, that the previously divergent term appearing at the right-hand-side of (61) is now regularized to $i\mathop{}\!\mathrm{d}\chi(2\pi\epsilon)^{-1/2}\partial_{\chi}\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$. As expected, the smaller the intrinsic uncertainty $\epsilon$ associated with the clock measurement is, the larger the quantum fluctuations on the clock momentum must be.
Note that if, instead of $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$, it is the effect density $\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ that is used to define the scalar field operator $\hat{\chi}_{\epsilon}$ and the relational observables $\hat{O}_{\chi;\epsilon}$ in (28) and (39) respectively, then the results (3.2) and (45) continue to hold to lowest order in $\epsilon$. In particular, when evaluating this new definition of relational observables to coherent states, up to corrections of order $\epsilon$,
@@
showing that (to leading order) the density $\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ produces the same number and volume effective relational observables defined for coherent peaked states [76]. Note, however, that this family of coherent peaked states allow small quantum fluctuations (in appropriate regimes) on both the clock observable and its momentum by including a non-trivial phase in the peaking function (58) [76]. An important difference is that the perspective adopted here is state-independent, as the quantum measurement of the clock is captured through the density $\hat{\mathcal{E}}_{\chi;\epsilon}(\mathop{}\!\mathrm{d}\chi)$ only, while (when appropriate) semi-classicality requirements should be imposed on states.